OpenStudy (anonymous):

hat is 8 with the exponent of -1/3 simplify to

OpenStudy (anonymous):

1/2

OpenStudy (diyadiya):

\[\Huge 8^ \frac{-1}{3} = \frac{1}{8^ \frac{1}{3}} = \frac{1}{\sqrt[3]{8}}\]